If 5x-2y=7 and xy=2, find the value of  5x+2y whole sqaure

Grade:8

1 Answers

Carmel Emmanuel
8 years ago
Substitute value of either x or y in xy=2 
Let substitute x 
x=2/y
5(2/y) -2y=7
U get a quadratic equation using quadratic formula u get a positive root of y=1(approx as it was comin something 1.05 by rounding off)
Now u get x=2
Now [5(2)+2(1)]2 = 144